Below is the time-phased baseline budget for a project and a table showing the progress as of End of Period 7.

Baseline PV

Task

Total PV

Time Period

0

1

2

3

4

5

6

7

8

9

10

11

A

30

15

15

B

40

20

20

C

80

10

10

10

10

20

20

D

30

5

5

5

5

10

E

20

5

5

5

5

F

40

10

20

5

5

G

10

5

5

PV by Period

15

15

20

35

15

20

30

55

30

5

5

5

Cumulative

15

30

50

85

100

120

150

205

235

240

245

250

As the end of Period 7, the project has made the following progress and incurred the following costs:

Task

% Complete

AC

A

100%

35

B

100%

45

C

40%

35

D

60%

20

E

50%

15

F

30%

10

Total

160

Question C13Q1:

What is the schedule variance? (Type a value rounded to whole dollars)

SV =

Question C13Q2:

What is the cost variance? (Type a value rounded to whole dollars)

CV =

Question C13Q3:

What fraction of the project has been completed in terms of original budget? Type a decimal below (rounded to 3 decimals. For example, type 0.456; DO NOT type 45.6%)

Project completion fraction (PICB):

Question C13Q4:

Forecast the total costs at completion based on current project cost and progress information. Type a value below (rounded to whole dollars while dropping the $ sign).

Forecasted total costs at completion (EAC):

Question C13Q5:

To complete the project within the original budget, what would be the required cost efficiency (TCPI)? Type a decimal value below (rounded to two decimals).

TCPI:

Question C13Q6:  Calculate the projected cost variance when the project is complete (rounded to whole dollars while dropping the $ sign).

Cost Variance (VAC):

36 points   
Answer #1

Total planned value at the end of week 7 = 205

earned value at the end of week 7 = 100% of A+100% of B + 40% of C +60% of D +50% of E +30% of F

= 30+40+32+18+10+12 =142

SV = EV -PV = 142-205 = -63

CV = EV-AC = 142-160 =-18

Fraction of project completed in terms of budget = EV / BAC = 142 /250 = 0.568

EAC = actual costs (AC) + budget at completion (BAC) – earned value (EV)

= 160+250-142 = 268
Note: The current Chegg policy permits me to answer only first four parts of a problem, Inconvenience is regretted.